Use elimination to solve: 6x-y/4=-9
3x+2y=-30

Answer by ankor@dixie-net.com(22740)

Use elimination to solve:
6x - y/4 = -9
3x + 2y = -30
:
Get rid of the denominator in the 1st equation by multiplying it by 4
4(6x) - y = 4(-9)
24x - y = -36
Now multiply this equation by 2 and add to the 2nd equation:
:
48x - 2y = -72
3x + 2y = -30
-----------------adding eliminates y; find x
51x + 0y = -102
x = -102/51
x = -2
:
Find y using the 2nd equation:
3(-2) + 2y = -30
-6 + 2y = -30
2y = -30 + 6
2y = -24
y = -24/2
y = -12
:
:
Check solutions using the 1st equation
6x - y/4 = - 9
6(-2) - (-12/4) = -9
-12 + 3 = -9; minus a minus is a plus